OpenStudy (anonymous):

Solve for x and y for the system: 4x - 6y = -20 3x + 6y = 48 What is x + y?

OpenStudy (anonymous):

We'll do this by substitution (Gaussian Elimination isn't worth it for 2 equations).

OpenStudy (kaederfds):

4x-6y=-20 3x+6y=48 --------- + 7x=28 x=4 3.4+6y=48 12+6y=48 6y=36 y=6 x+y=10

OpenStudy (anonymous):

\[\begin{align} 4x=6y-20\\ x=\tfrac{3}{2}y-5\\ \tfrac{9}{2}y-15+6y=48\\ 21y=126\\ y=6\\ x=9-5=4\\ x+y=10 \end{align}\]

OpenStudy (anonymous):

Thank you so very much- I had 6.. :(